# Traversal Verification for Speculative Tree Decoding

### v1

@@ Abstract
Speculative decoding is a promising approach for accelerating large language models. The primary idea is to use a lightweight draft model to speculate the output of the target model for multiple subsequent timesteps, and then verify them in parallel to determine whether the drafted tokens should be accepted or rejected. To enhance acceptance rates, existing frameworks typically construct token trees containing multiple candidates in each timestep. However, their reliance on token-level verification mechanisms introduces two critical limitations: First, the probability distribution of a sequence differs from that of individual tokens, leading to suboptimal acceptance length. Second, current verification schemes begin from the root node and proceed layer by layer in a top-down manner. Once a parent node is rejected, all its child nodes should be discarded, resulting in inefficient utilization of speculative candidates. This paper introduces Traversal Verification, a novel speculative decoding algorithm that fundamentally rethinks the verification paradigm through leaf-to-root traversal. Our approach considers the acceptance of the entire token sequence from the current node to the root, and preserves potentially valid subsequences that would be prematurely discarded by existing methods. We theoretically prove that the probability distribution obtained through Traversal Verification is identical to that of the target model, guaranteeing lossless inference while achieving substantial acceleration gains. Experimental results on different large language models and multiple tasks show that our method consistently improves acceptance length and throughput over existing methods.
@@ 1 Introduction
Large Language Models (LLMs) have been widely adopted due to their exceptional performance across various natural language processing tasks [9, 24, 33] . However, the massive parameters and the autoregressive generation scheme of transformer decoder-only [29] LLMs limit the generation speed. Speculative decoding [18, 3] is an acceleration technique which employs a lightweight model (draft model) with fewer parameters to speculate the output tokens of the original LLM (target model) for several future timesteps, then feed the drafted tokens into the target model in parallel. After getting the probability distribution of the target model, speculative decoding sequentially determines the acceptance or rejection of each token based on their probabilities in both target and draft models. If a token is rejected, a new token will be resampled and all subsequent tokens should be discarded.
To further improve acceleration performance, existing methods [22, 4, 20, 13, 34] generate multiple candidates at each drafting timestep, forming a tree of drafted tokens. However, these methods generally inherit the token-level verification mechanism from vanilla speculative decoding to tree scenarios, resulting in suboptimal acceptance lengths in tree decoding. To be more specific, firstly, the probability distribution of a token sequence differs from that of an individual token. Vanilla speculative decoding determines acceptance based on per-token probabilities, which sacrifices global optimality for sequence-level acceptance. Secondly, existing tree decoding methods start verification from the root node of the tree, and proceed layer by layer in a top-down manner. Once a parent node is rejected, all its child nodes will be discarded accordingly, resulting in the wasting of drafted tokens.
@@
We conducted experiments on Llama3 [9] series and Llama2 [28] using various tree structures. The experiments were performed on the Spec-Bench dataset [31], which encompasses six different tasks: multi-turn conversation, translation, summarization, question answering, mathematical reasoning, and retrieval-augmented generation. Experimental results demonstrate that Traversal Verification consistently outperforms existing decoding methods by 2.2%-5.7% in average acceptance length across diverse tasks with different tree architectures. Additionally, Traversal Verification could potentially achieve greater improvements for deeper and larger decoding trees.
@@ 2.1 Speculative Decoding
Speculative decoding, also known as speculative sampling [18, 3], is a lossless LLM acceleration algorithm. In speculative decoding, a draft model first generates a chain of $\gamma$ new tokens (i.e., one token per timestep for the next $\gamma$ timesteps), then the drafted tokens are fed into the target model in parallel to get the target distribution.
@@
According to Algorithm 1, if token $b$ is sampled, it will be accepted directly because $\mathcal{M}_{b}(b)>\mathcal{M}_{s}(b)$. Similarly, $c$ will be accepted if sampled. If $a$ is sampled, the acceptance probability is $\frac{\mathcal{M}_{b}(a)}{\mathcal{M}_{s}(a)}=0.5$. Thus, the probability of generating token $a$ is (sample $\mathbb{P}(\text{sample }a)\times\mathbb{P}(\text{accept }a)=0.3$, which is equal to $\mathcal{M}_{b}(a)$. These cases correspond to the diagonal entries in Table 1, highlighted in green.
@@ 2.2 Recursive Rejection Sampling
Recursive Rejection Sampling (RRS) samples multiple candidates at each timestep and recursively verifies them, as described in Algorithm 2. Recent works [4, 20, 13, 34] further refine RRS into RRS without replacement (RRSw), where the probability of a rejected token in $\mathcal{M}_{s}$ is set to zero, and then normalize $\mathcal{M}_{s}$ of the remaining candidates. RRSw prevents repeated sampling and rejection of the same token, especially for low-temperature situations, thereby improving overall acceptance rates.
@@ 3.1 Traversal Verification
    Bottom-up verification. Traversal Verification generally operates in a bottom-up manner, starting verification from leaf nodes (i.e., deeper layers) and progressing toward the root, while token-level tree decoding follows a top-down approach, verifying nodes layer by layer from shallow to deep.
@@
    Sequence-level acceptance. Traversal Verification incorporates the joint probability distribution of the token sequence, rather than relying solely on per-token probabilities. Once a token is accepted, the entire sequence from the current node to root is accepted.
@@ 3.2 An Intuitive Example of Traversal Verification
We now demonstrate Traversal Verification using the same illustrative case as introduced in Section 2.2. Following Algorithm 3, for the tree structure in Figure 1, the traversal order is $X_{3}\rightarrow X_{4}\rightarrow X_{1}\rightarrow X_{5}\rightarrow X_{2}$. Consider a tree with nodes sampled as $[X_{1},X_{2},X_{3},X_{4},X_{5}]=[a,c,b,c,a]$ as an intuitive example. We present the detailed process of Traversal Verification in Figure 2.
@@
When $X_{1}X_{3}$ is rejected, we delete the last node $X_{3}$ and then the first chain becomes $X_{1}X_{4}$. According to Line 11-13 in Algorithm 3, since $[p(X_{1})\mathcal{M}_{b}(a)-\mathcal{M}_{s}(a)]_{+}=0$ and $[p(X_{1})\mathcal{M}_{b}(c)-\mathcal{M}_{s}(c)]_{+}=0.05$, the new $p^{\prime}(X_{1})$ and the acceptance rate of chain $X_{1}X_{4}$ are updated as
@@
(accept $\mathbb{P}(\text{accept }X_{1}X_{4})=\min\left(p^{\prime}(X_{1})\cdot\frac{\mathcal{M}^{\prime}_{b}(X_{4})}{\mathcal{M}^{\prime}_{s}(X_{4})},1\right)\approx 0.637.$
@@ 3.3 Theoretical Guarantees
We first formally define the stochastic sampling tree as follows:
Definition 3.1 (Stochastic sampling tree based on $\mathcal{M}_{s}$).
Let $T$ be a tree and ${\rm Path}(v)$ represent the list of child indices denoting the path of a chain $v\subset T$. For example, if $v$ is the chain from the root to the $3^{rd}$ child of the root’s $2^{nd}$ child, then ${\rm Path}(v)=[2,3]$. The tree $T$ is called to be a valid sampling tree on distribution $\mathcal{M}_{s}$, if for any $\mathrm{Path}(v)$ of $T$ with $\gamma_{v}$ being the length of ${\rm Path}(v)$, the induced probability measure satisfies:
$\mathbb{P}(v=(X_{0},X_{1},\dots,X_{i}))=\prod_{k=1}^{i}\mathcal{M}_{s}(X_{k}|X^{k-1}),\quad\forall 1\leqslant i\leqslant\gamma_{v},$
where $X_{0}$ is the root and $\mathbb{P}(v=X_{0})=\mathcal{M}_{s}(X_{0})=1$.
Given a stochastic sampling tree $T$, we prove that Traversal Verification serves as a valid tree verification algorithm. A valid tree verification algorithm is defined as follows:
@@
Let $T$ be a valid sampling tree on $\mathcal{M}_{s}$ with root node $X_{0}$. For all chains $v=(X_{0},\dots,X_{\gamma_{v}})\subset T$ where $\gamma_{v}$ denotes the depth of chain $v$, a tree verification algorithm $\mathcal{A}_{\mathrm{ver}}$ takes the tree $T$, draft model distributions $\mathcal{M}_{s}(\cdot|X^{i})$, $\forall i<\gamma_{v}$ and target model distributions $\mathcal{M}_{b}(\cdot|X^{i})$, $\forall i\leqslant\gamma_{v}$ as inputs, and outputs an accept chain $X^{\tau}=(X_{0},\dots,X_{\tau})\subset T$ where $\tau\leqslant\max_{v\subset T}\gamma_{v}$ and an additional token $Y$.
@@
For example, SpecInfer [22, Theorem 4.2] is a valid tree verification algorithm. In the case where the sampling tree degenerates into a single chain, both the vanilla token verification [18, Appendix.A] and Block Verification [26, Theorem 1] are valid chain verifications.
@@
When the sampling tree reduces to one single chain, for any valid chain verification algorithm VERIFY in Definition 3.2, let $N_{\rm traversal}$, $N_{\rm block}$ and $N_{\rm verify}$ be the number of accepted tokens in Traversal Verification, Block Verification [26] and VERIFY, respectively, then for any given distributions $\mathcal{M}_{s},\mathcal{M}_{b}$ and draft chain $T$, we have
@@
Theoretical foundations and design motivation of Traversal Verification. The core idea of proving the losslessness (Theorem 3.3) of Traversal Verification lies in exploiting its self-similarity. The self-similarity of Traversal Verification implies that, for any parent node $A$ in the given sampling tree $T$, before determining the acceptance of $A$, all its descendant nodes have already been processed through the same traversal mechanism. In other words, every local subtree within the sampling tree $T$ essentially operates as a scaled-down instance of the Traversal Verification mechanism. Consequently, we can employ mathematical induction on the number of descendant nodes to establish the critical Lemma A.2, from which Theorem 3.3 (the lossless theorem) directly follows as a corollary.
For the single-chain optimality of Traversal Verification (Theorem 3.4), the key proof idea is to ensure that Traversal Verification achieves the highest possible acceptance probability at each node, aligning with Block Verification. Assume that the highest acceptance rate for any parent node $A$ is $p(A)$. As a bottom-up verification framework, the target probability distribution for child nodes of $A$ should be $p(A)\mathcal{M}_{b}$. By introducing a pseudo-child node with target probability $(1-p(A))$, we can apply RRSw to transport the draft distribution $\mathcal{M}_{s}$ to the target distribution $p(A)\mathcal{M}_{b}$ combining with $(1-p(A))$. This motivation directly leads to the formulations (1)–(3) of Traversal Verification. Since Block Verification is an optimal valid chain verification algorithm [26, Theorem 2], Traversal Verification inherits this optimality in the single-chain case (see Theorem 3.4).
@@ 4.1 Experimental Setup
We mainly conduct experiments on the Llama3 [9] series, using Llama3.2-1B-Instruct as the draft model and Llama3.1-8B-Instruct as the target model. We also include Llama-68M [22] with Llama2-7b [28] as the draft and target model, which is widely adopted in existing speculative decoding researches [4, 11, 12, 25].
@@
We perform experiments on the Spec-Bench dataset [31], which includes 80 instances from each of six distinct domains: multi-turn conversation (MT-Bench [35]), translation (WMT14 DE-EN [1]), summarization (CNN/Daily Mail [23]), question answering (Natural Questions [17]), Mathematical reasoning (GSM8K [5]), retrieval-augmented generation (DPR [15]).
@@
We evaluate the performance of our method using two metrics: acceptance length and token generation speed. Acceptance length is the number of tokens generated per drafting-verification cycle, which reflects the theoretical performance of the verification method. We also include the actual throughput for a comprehensive comparison. Notably, since Traversal Verification is a lossless speculative decoding method, there is no need to evaluate the generation quality.
@@
For token-level tree verification, we adopt the RRSw implementation in EAGLE [20] from Spec-Bench [31] open source repository. All experiments are conducted on a single NVIDIA RTX A6000 GPU with PyTorch backend. Due to inherent randomness in sampling, we conduct three independent runs for each case and report the average as the result.

### v2

@@ Abstract
Speculative decoding is a promising approach for accelerating large language models. The primary idea is to use a lightweight draft model to speculate the output of the target model for multiple subsequent timesteps, and then verify them in parallel to determine whether the drafted tokens should be accepted or rejected. To enhance acceptance rates, existing frameworks typically construct token trees containing multiple candidates in each timestep. However, their reliance on token-level verification mechanisms introduces two critical limitations: First, the probability distribution of a sequence differs from that of individual tokens, leading to suboptimal acceptance length. Second, current verification schemes begin from the root node and proceed layer by layer in a top-down manner. Once a parent node is rejected, all its child nodes should be discarded, resulting in inefficient utilization of speculative candidates. This paper introduces Traversal Verification, a novel speculative decoding algorithm that fundamentally rethinks the verification paradigm through leaf-to-root traversal. Our approach considers the acceptance of the entire token sequence from the current node to the root, and preserves potentially valid subsequences that would be prematurely discarded by existing methods. We theoretically prove that the probability distribution obtained through Traversal Verification is identical to that of the target model, guaranteeing lossless inference while achieving substantial acceleration gains. Experimental results on various models and multiple tasks demonstrate that our method consistently improves acceptance length and throughput over token-level verification.
@@ 1 Introduction
Large Language Models (LLMs) have been widely adopted due to their exceptional performance across various natural language processing tasks [grattafiori2024llama3herdmodels, DBLP:journals/corr/abs-2303-08774, qwen2] . However, the massive parameters and the autoregressive generation scheme of transformer decoder-only [DBLP:conf/nips/VaswaniSPUJGKP17] LLMs limit the generation speed. Speculative decoding [DBLP:conf/icml/LeviathanKM23, DBLP:journals/corr/abs-2302-01318] is an lossless acceleration technique which employs a lightweight model (draft model) with fewer parameters to speculate the output tokens of the original LLM (target model) for several future timesteps, then feed the drafted tokens into the target model in parallel. After getting the probability distribution of the target model, speculative decoding determines the acceptance or rejection of each token based on their probabilities in both target and draft models. If a token is rejected, a new token will be resampled and all subsequent tokens should be discarded.
To further improve acceleration performance, existing methods [DBLP:conf/asplos/MiaoOZCWZWZYSSC24, ChenSequoia, LiEAGLE, JeonRSD, YangMCSD] generate multiple candidates at each drafting timestep, forming a tree of drafted tokens. However, these methods generally inherit the token-level verification mechanism from vanilla speculative decoding to tree scenarios, resulting in suboptimal acceptance lengths in tree decoding. To be more specific, firstly, the probability distribution of a token sequence differs from that of an individual token. Vanilla speculative decoding determines acceptance based on per-token probabilities, which sacrifices global optimality for sequence-level acceptance. Secondly, existing tree decoding methods start verification from the root node of the tree, and proceed layer by layer in a top-down manner. Once a parent node is rejected, all its child nodes will be discarded accordingly, resulting in the wasting of drafted tokens.
@@
We conducted experiments on Llama3 [grattafiori2024llama3herdmodels] series and Llama2 [llama2] using various tree structures. The experiments were performed on the Spec-Bench dataset [XiaSpecbench], which encompasses six different tasks: multi-turn conversation, translation, summarization, question answering, mathematical reasoning, and retrieval-augmented generation. Experimental results demonstrate that Traversal Verification consistently outperforms existing decoding methods by 2.2%-5.7% in average acceptance length across diverse tasks with different tree architectures. Additionally, Traversal Verification could potentially achieve greater improvements for deeper and larger decoding trees.
@@ 2.1 Speculative Decoding
Speculative decoding, also known as speculative sampling [DBLP:conf/icml/LeviathanKM23, DBLP:journals/corr/abs-2302-01318], is a lossless LLM acceleration algorithm. In speculative decoding, a draft model first generates a chain of $\gamma$ new tokens (i.e., one token per timestep for the next $\gamma$ timesteps), then the drafted tokens are fed into the target model in parallel to get the target distribution.
@@
According to Algorithm 1, if token $b$ is sampled, it will be accepted directly because $\mathcal{M}_{b}(b)>\mathcal{M}_{s}(b)$. Similarly, $c$ will be accepted if sampled. If $a$ is sampled, the acceptance probability is $\frac{\mathcal{M}_{b}(a)}{\mathcal{M}_{s}(a)}=0.5$. Thus, the probability of generating token $a$ is $\mathbb{P}(\text{sample }a)\times\mathbb{P}(\text{accept }a)=0.3$, which is equal to $\mathcal{M}_{b}(a)$. These cases correspond to the diagonal entries in Table 1, highlighted in green.
@@ 2.2 Recursive Rejection Sampling
Recursive Rejection Sampling (RRS) samples multiple candidates at each timestep and recursively verifies them, as described in Algorithm 2. Recent works [ChenSequoia, LiEAGLE, JeonRSD, YangMCSD] further refine RRS into RRS without replacement (RRSw), where the probability of a rejected token in $\mathcal{M}_{s}$ is set to zero, and then normalize $\mathcal{M}_{s}$ of the remaining candidates. RRSw prevents repeated sampling and rejection of the same token, especially for low-temperature situations, thereby improving overall acceptance rates.
@@ 3.1 Traversal Verification
    Bottom-up verification. Traversal Verification generally operates in a bottom-up manner, starting verification from leaf nodes (i.e., deeper layers) and progressing toward the root, while token-level tree decoding follows a top-down approach, verifying nodes layer by layer from shallow to deep. Details about traversal order are provided in Appendix E.
@@
    Sequence-level acceptance. Traversal Verification incorporates the joint probability distribution of the token sequence, rather than relying solely on per-token probabilities. The acceptance rate at each node represents the sequence-level acceptance rate from the current node to the root. Thus, once a token is accepted, the entire sequence from the current node to root is accepted.
@@ 3.2 An Intuitive Example of Traversal Verification
We now demonstrate Traversal Verification using the same illustrative case as introduced in Section 2.2. Following Algorithm 3, for the tree structure in Figure 1, the traversal order is $X_{3}\rightarrow X_{4}\rightarrow X_{1}\rightarrow X_{5}\rightarrow X_{2}$. Consider a tree with nodes sampled as $X_{1},X_{2},X_{3},X_{4},X_{5}]=[a,c,b,c,a]$ as an intuitive example. We present the detailed process of Traversal Verification in Figure 2.
@@
When $X_{1}X_{3}$ is rejected, we delete the last node $X_{3}$ and then the first chain becomes $X_{1}X_{4}$. According to Line 11-13 in Algorithm 3, since $p(X_{1})\mathcal{M}_{b}(a)-\mathcal{M}_{s}(a)]_{+}=0$ and $p(X_{1})\mathcal{M}_{b}(c)-\mathcal{M}_{s}(c)]_{+}=0.05$, the new $p^{\prime}(X_{1})$ and the acceptance rate of chain $X_{1}X_{4}$ are updated as
@@
$\mathbb{P}(\text{accept }X_{1}X_{4})=\min\left(p^{\prime}(X_{1})\cdot\frac{\mathcal{M}^{\prime}_{b}(X_{4})}{\mathcal{M}^{\prime}_{s}(X_{4})},1\right)\approx 0.637.$
@@ 3.3 Theoretical Guarantees
We first formally define the decoding tree under autoregressive generation as follows:
Definition 3.1 (Decoding tree under autoregressive generation).
Let $\mathcal{M}_{s}$ be a given distribution and $T$ be a decoding tree rooted at $X_{0}$ under autoregressive generation. For all chains $v=(X_{0},\dots,X_{\gamma_{v}})\subset T$ where $\gamma_{v}$ denotes the depth of chain $v$, if all child nodes of $v$ are generated according to the conditional distribution $\mathcal{M}_{s}(\cdot|v)$ (with or without replacement), then the sampling tree $T$ is termed a decoding tree based on $\mathcal{M}_{s}$ under autoregressive generation.
For brevity, we hereafter refer to a tree satisfying the above definition as a decoding tree.
Given an decoding tree $T$, we prove that Traversal Verification serves as a valid tree verification algorithm. A valid tree verification algorithm is defined as follows:
@@
Let $T$ be a decoding tree defined as Definition 3.1. For all chains $v=(X_{0},\dots,X_{\gamma_{v}})\subset T$ with depth $\gamma_{v}$, a tree verification algorithm $\mathcal{A}_{\mathrm{ver}}$ takes the tree $T$, draft model distributions $\mathcal{M}_{s}(\cdot|X^{i})$, $\forall i<\gamma_{v}$ and target model distributions $\mathcal{M}_{b}(\cdot|X^{i})$, $\forall i\leqslant\gamma_{v}$ as inputs, and outputs an accept chain $X^{\tau}=(X_{0},\dots,X_{\tau})\subset T$ where $\tau\leqslant\max_{v\subset T}\gamma_{v}$ and an additional token $Y$.
@@
For example, SpecInfer [DBLP:conf/asplos/MiaoOZCWZWZYSSC24, Theorem 4.2] is a valid tree verification algorithm. In the case where the sampling tree degenerates into a single chain, both the vanilla token verification [DBLP:conf/icml/LeviathanKM23, Appendix.A] and Block Verification [blockV, Theorem 1] are valid chain verifications.
@@
When the sampling tree reduces to one single chain, for any valid chain verification algorithm VERIFY in Definition 3.2, let $N_{\rm traversal}$, $N_{\rm block}$ and $N_{\rm verify}$ be the number of accepted tokens in Traversal Verification, Block Verification [blockV] and VERIFY, respectively, then for any given distributions $\mathcal{M}_{s},\mathcal{M}_{b}$ and draft chain $T$, we have
@@
Discussions on theoretical foundations and design motivation of Traversal Verification. The core idea of proving the losslessness (Theorem 3.3) of Traversal Verification lies in exploiting its self-similarity. The self-similarity of Traversal Verification implies that, for any parent node $A$ in the given sampling tree $T$, before determining the acceptance of $A$, all its descendant nodes have already been processed through the same traversal mechanism. In other words, every local subtree within the sampling tree $T$ essentially operates as a scaled-down instance of the Traversal Verification mechanism. Consequently, we can employ mathematical induction on the number of descendant nodes to establish the critical Lemma A.2, from which Theorem 3.3 (the lossless theorem) directly follows as a corollary.
For the single-chain optimality of Traversal Verification (Theorem 3.4), the key proof idea is to ensure that Traversal Verification achieves the highest possible acceptance probability at each node, aligning with Block Verification. Assume that the acceptance rate for a parent node $A$ is $P(A)$. As a bottom-up verification framework, the target probability distribution for child nodes of $A$ should be $P(A)\mathcal{M}_{b}$. By introducing a pseudo-child node with target probability (1P(A))(1-P(A)), we can apply RRSw to transport the draft distribution $\mathcal{M}_{s}$ to the target distribution $P(A)\mathcal{M}_{b}$ combining with (1P(A))(1-P(A)). We refer to the above process as the sequence-level RRSw method. Comprehensive details are provided in Appendix F. This motivation directly leads to the formulations (1)–(3) of Traversal Verification. Since Block Verification is an optimal valid chain verification algorithm [blockV, Theorem 2], Traversal Verification inherits this optimality in the single-chain case (see Theorem 3.4).
@@ 4.1 Experimental Setup
We mainly conduct experiments on the Llama3 [grattafiori2024llama3herdmodels] series, using Llama3.2-1B-Instruct as the draft model and Llama3.1-8B-Instruct as the target model. We also include Llama-68M [DBLP:conf/asplos/MiaoOZCWZWZYSSC24] with Llama2-7b [llama2] as the draft and target model, which is widely adopted in existing speculative decoding researches [ChenSequoia, Asps, TowardsOptimal, SpecHub].
@@
We perform experiments on the Spec-Bench dataset [XiaSpecbench], which includes 80 instances from each of six distinct domains: multi-turn conversation (MT-Bench [mt-bench]), translation (WMT14 DE-EN [wmt14]), summarization (CNN/Daily Mail [cnndm]), question answering (Natural Questions [natural_questions]), Mathematical reasoning (GSM8K [gsm8k]), retrieval-augmented generation (DPR [DPR]).
@@
We evaluate the performance of our method using two metrics: acceptance length and token generation speed. Acceptance length is the number of tokens generated per drafting-verification cycle, which reflects the theoretical performance of the verification method. We also include the actual throughput for a comprehensive comparison. It is worth noting that there may be slight variations in acceptance length according to differences in statistical methods, and we provide detailed discussions and additional experimental results on this issue in Appendix D.
@@
For token-level tree verification, we adopt the RRSw implementation in EAGLE [LiEAGLE] from Spec-Bench [XiaSpecbench] open source repository. All experiments are conducted on a single NVIDIA RTX A6000 GPU with PyTorch backend. Due to inherent randomness in sampling, we conduct three independent runs for each case and report the average as the result.
Measurement of Generation Quality.
Traversal Verification is theoretically a lossless speculative decoding technique, which suggests that evaluating its generation quality should not be mandatory. However, recognizing that some readers may seek assurance regarding this guarantee, we present the measurements of generation quality as a supporting reference for losslessness. Please consult Appendix C for the detailed experimental findings.